Fermi National Laboratory


Accelerator Update

Wednesday January 28
The day shift began with Operations monitoring store 3197, with the Antiproton Source (Pbar) stacking, and with MiniBooNE taking beam.

At 8:21 AM, the Main Injector (MI) stopped accelerating beam. Operators contacted experts. Recycler experts noted a problem with their machine also. RF experts rebooted the low level RF for MI at 8:48 AM and beam resumed.

Operators began preparing MI, Booster, and Linac for an access at 10:04 AM. This access halted stacking and MiniBooNE beam.

Operations established beam to the Linac dump at 3:39 PM.

Pbar experts stopped their studies and resumed stacking 10:24 PM.

Operations terminated store 3197 at 10:54 PM.

At 11:10 PM, Operations put the TeV through a dry squeeze.

Thursday January 29
The midnight shift began with Tevatron (TeV) experts conducting orbit smoothing, and with Pbar experts conducting studies.

At 12:39 AM, a MiniBooNE correction magnet power supply (E:VT865) tripped off. Operators managed to get the supply back on by 12:43 AM. It tripped again at 1 AM. Operators reset it.

Due to an unfortunate side effect of Pbar studies, 100 mA of antiprotons were lost from the stack at 2:57 AM.

Pbar resumed stacking at 3:09 AM.

At 6 AM and again at 6:34 AM, Operations established 12x0 study-stores.

Operations put the TeV through a dry squeeze at 8:47 AM.

Operations put the TeV into shot setup at 10 AM and established store 3206 at 12:15 PM. Its initial luminosity was 32.87E30. Pbar resumed stacking.

Recycler experts took protons into their machine at 1:21 PM.

At 1:30 PM, a Booster kicker (MKS03) began sparking. Technicians discovered a bad resistor. They had it replaced and the kicker working by 3:15 PM.

Friday January 30
The midnight shift began with Operations monitoring store 3206, with MiniBooNE taking beam, and with Pbar stacking.

A Booster power supply (ORBMP) stopped putting out current at 12:09 AM. Operators had it working normally within five minutes.

At 2:34 AM, the TeV suffered a sector F4 quench.

Operations made D-zero (D0) ready for an access at 6:34 AM. During the night the D0 experimenters saw an alarm that may have been due to a water leak in their collision hall. At 7:41 AM, they reported that they had to open their detector to make repairs. They estimated that this would take about one shift, eight hours.

Today's Plans
The plan for today is to allow the D0 experimenters to repair their detector. TeV maintenance will also occur during this downtime.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 1/30/2004   email Fermilab